# Rethinking Domain-Specific LLM Benchmark Construction: A Comprehensiveness-Compactness Approach

### v2

@@ Abstract
Numerous benchmarks have been built to evaluate the domain-specific abilities of large language models (LLMs), highlighting the need for effective and efficient benchmark construction. Existing domain-specific benchmarks primarily focus on the scaling law, relying on massive corpora for supervised fine-tuning or generating extensive question sets for broad coverage. However, the impact of corpus and question-answer (QA) set design on the precision and recall of domain-specific LLMs remains unexplored. In this paper, we address this gap and demonstrate that the scaling law is not always the optimal principle for benchmark construction in specific domains. Instead, we propose Comp-Comp, an iterative benchmarking framework based on a comprehensiveness-compactness principle. Here, comprehensiveness ensures semantic recall of the domain while compactness enhances precision, guiding both corpus and QA set construction. To validate our framework, we conducted a case study in a well-renowned university, resulting in the creation of XUBench, a large-scale and comprehensive closed-domain benchmark. Although we use the academic domain as the case in this work, our Comp-Comp framework is designed to be extensible beyond academia, providing valuable insights for benchmark construction across various domains.
Rethinking Domain-Specific LLM Benchmark Construction:
A Comprehensiveness-Compactness Approach
Rubing Chen1, Jiaxin Wu1, Jian Wang1, Xulu Zhang1, Wenqi Fan1, Chenghua Lin2, Xiao-Yong Wei1,*, Qing Li1 1The Hong Kong Polytechnic University  2The University of Manchester
@@ 1 Introduction
Figure 1: Illustration of comprehensiveness and compactness in benchmarking. (a) Comprehensiveness ensures broad semantic coverage of the domain, while compactness improves precision, leading to a more uniform data distribution in domain corpora. (b) Empirical semantic distributions of five representative benchmarks in the medical domain.
@@
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks OpenAI (2024). While general-purpose LLMs excel in broad applications, domain-specific LLMs have emerged as critical tools for delivering precise and accurate responses tailored to specific user needs Fei et al. (2023); Cai et al. (2024); Chang et al. (2024). These models are particularly valuable in fields where fixed-answer questions are prevalent, such as law, medicine, and academia, where precision is paramount Ge et al. (2024). However, the development of domain-specific LLMs necessitates the creation of domain-specific benchmarks that can effectively evaluate their performance. Such benchmarks must balance precision—ensuring accurate responses within the domain—with recall—maintaining broad coverage of the domain’s semantic space to avoid catastrophic forgetting, where models lose proficiency in broader contexts Luo et al. (2023). This dual requirement underscores the need for effective and efficient approaches to benchmark construction.
Previous studies for constructing domain-specific benchmarks primarily rely on traditional scaling laws, leveraging extensive corpora or large question sets to achieve broad coverage Cai et al. (2024); Fei et al. (2023). While this approach is common, it is not without limitations. Many benchmarks are built using data from skill examinations, practice exercises, or expert-constructed datasets, often adhering to scaling laws without critical evaluation of their effectiveness Liu et al. (2023); Ling et al. (2023). Others, despite covering multiple question types, fail to systematically analyze the semantic alignment between questions and the underlying corpus Yue et al. (2023); Cui et al. (2023). Additionally, some benchmarks employ LLMs for scoring but overlook the consistency between training data and evaluation sets Zhang et al. (2023); Yang et al. (2024). These shortcomings raise critical questions: What constitutes a sufficient benchmark size for a specific domain? How do training corpora and benchmark coverage relate to model performance? How many questions are necessary to assess a model effectively and efficiently? Addressing these gaps requires a more balanced approach to benchmark design, one that considers both horizontal coverage and vertical depth.
@@
In this paper, we propose Comprehensiveness-Compactness (Comp-Comp), a novel framework for constructing domain-specific benchmarks that dynamically balances the semantic distribution of corpora and benchmark questions. As illustrated in Figure 1, Comp-Comp iteratively expands the range of corpora and questions, continuously assessing their comprehensiveness (maximizing recall) and compactness (optimizing precision). The framework introduces three crucial advantages: (1) Comprehensiveness and Compactness Monitoring: We encode the semantics of corpora and questions into a unified semantic space, using Gaussian Kernel Density Estimation (KDE) to measure and iteratively improve these metrics Terrell and Scott (1992). (2) Corpora-Question Interaction: We ensure comprehensive evaluation by generating new questions from underrepresented areas in the corpora. (3) Superiority in Closed Domains: We prioritize achieving a balance between comprehensiveness and compactness, ensuring the framework excels in constructing closed-domain benchmarks that demand precise corpus evaluation and accurate answer verification.
@@
To validate our framework, we conducted a case study in a well-renowned university and applied Comp-Comp to develop XUBench, one of the first and most extensive closed-domain benchmarks for the academic domain (see Table 1). XUBench includes 24.9k questions across various types and incorporates user-interest-oriented forum discussions, setting a new standard for domain-specific benchmarking. This case study demonstrates the practicality and effectiveness of our approach, providing a scalable model for future benchmark construction in other domains.
@@
In summary, our contributions are as follows: (1) We propose the Comp-Comp framework, which introduces a principled approach to balancing comprehensiveness and compactness in domain-specific benchmark construction. (2) We develop XUBench, a large-scale, closed-domain benchmark for academia, which addresses the gap of lacking well-established academic LLM benchmarks. (3) Our method is extensible across domains, offering valuable insights for future benchmark construction. By addressing the limitations of existing methods and providing a systematic benchmark design, our work advances the field of domain-specific LLM evaluation and development.
@@ 2 Related Work
Domain-specific evaluation has seen significant advances across critical verticals, each adopting distinct strategies reflective of their unique demands. In the legal domain, systems like DISC-LawLLM Yue et al. (2023) employ a dual evaluation paradigm: standardized law examinations (MCQs with tiered difficulty) paired with ChatGPT, arbitrated subjective assessments. LawBench Fei et al. (2023) further refines this by grounding evaluations in Bloom’s Taxonomy Krathwohl (2002), systematically testing knowledge memorization, understanding, and application within Chinese legal contexts. The medical domain presents alternative approaches: ChatCounselor Liu et al. (2023) evaluates seven mental consultation dimensions via GPT-4, while CMtMedQA Yang et al. (2024) uses dialogue-based capacity assessments scored across three dimensions. Even within single domains, fragmentation persists: Figure 1(b) reveals disjointed semantic distributions across medical benchmarks, including MedExQA Kim et al. (2024), MMLU Hendrycks et al. (2021), PubMedQA Jin et al. (2019), and MedQA Jin et al. (2021), with clear boundaries between corpora despite shared topical grounding. Educational benchmarks like C-Eval Huang et al. (2024), though comprehensive, exhibit similar fragmentation, as noted in EduChat’s evaluation Dan et al. (2023). These efforts collectively highlight a critical gap: the absence of unified metrics to assess whether benchmarks sufficiently cover a domain’s semantic space or align with training corpora.
@@ 3.1 Comprehensiveness and Compactness Assessment
Given a set of data point $X=\{x_{1},x_{2},...,x_{m}\}$, its comprehensiveness is measured by computing its coverage of a space $S$ with data points $\{s_{1},s_{2},...,s_{n}\}$. To achieve this, we encode all the data points in $X$ and $S$ to the same semantic space as follows:
@@
On the other hand, the compactness of the corpora and QA is controlled by whether the inclusion of a new set of data points $X$ brings much semantic residency to the existing set $Y$. To this end, we measure the Pearson correlation coefficient Cohen et al. (2009) between the semantic distributions of $X$ and $Y$ as:
@@ 3.3 Mining User Interests for QA Generation
Moreover, in the QA generation, we propose that we should not only generate QAs based on the fact on the corpora $C$ but also should include user-interested questions. For example, the frequently asked questions and the questions discussed in public forums (e.g., Quora and Zhihu To this end, we add user-interested questions as a special gap points to $A_{gap}^{\prime}(C,Q)$.
4 XUBench: An Academic Benchmark
In this section, we employ the proposed framework to construct XUBench, a comprehensive and compact benchmark in the academic domain.
@@ 4.1 Assessment Guided Corpora Collection
Benchmark Domain Closeness QA Type # Questions
SciEval Cai et al. (2024) Science Open MCQ, QA 16,522
DISC-Law-Eval Yue et al. (2023) Law Open MCQ, MAQ, QA 2,863
Counseling Bench Liu et al. (2023) Medical Open QA 229
CMtMedQA Yang et al. (2024) Medical Open QA 70,000
COVID-QA Möller et al. (2020) Healthcare Closed QA 2,019
WC2014QA Zhang et al. (2016) Sports Closed QA 10,211
XUBench (Ours) Academic Closed Binary, MAQ, QA, MCQ 24,883
Table 1: Comparison between domain-specific benchmarks.
@@
Table 2: Statistics of the XUBench.
@@ 5 Experiments
In this section, we mainly explore how in-context learning (ICL) and supervised fine-tuning (SFT), which are two main solutions for domain-specific LLMs, perform on the XUBench and provide insights for the community.
@@
Table 3: The performance of In-Context Learning (ICL) and Supervised Fine-tuning (SFT) approaches (i.e., few-shot learning and RAG) on the XUBench with various types of questions.
@@ 5.1 Experimental Settings
We include both frozen and SFT LLMs with ICL prompts as baselines. For the ICL, we apply few-shot learning Parnami and Lee (2022) and Retrieval Augmented Generation (RAG) Lewis et al. (2020); Gao et al. (2023) to include domain knowledge in the prompt for LLMs. For the SFT, we use LoRA Hu et al. (2021) to fine-tune several LLM backbones. For LLM backones, we apply both close-sourced and open-sourced LLMs including GPT3.5, GPT4 OpenAI (2024), Vicuna-7B, Vicuna-13B Chiang et al. (2023), LLaMA3-8B Touvron et al. (2023), GLM4-9B GLM et al. (2024), and Qwen2-7B Bai et al. (2023).
@@
For the XUBench construction, we use $t_{c}=0.05$ in Eq. (7) for the corpora collection and $h=5.0$ in Eq.(3) and Eq. (4), and $t_{d}=0.6$ in Eq. (8) for the QA generation. We use Gte-Large Li et al. (2023) encoder to encode data chunks.
5.2 In-Context Learning on XUBench
Table 3 compares the performances of different in-context-learning techniques, i.e., few-shot learning and RAG, on answering several different QA sets of XUBench. Experiment results on more LLMs are listed in Appendix A’s Table 5. Based on the results, we have several findings:
@@ 5.2 In-Context Learning on XUBench
Our analysis demonstrates that RAG substantially improves precision and recall over vanilla LLMs and few-shot learning in domain-specific QA tasks. This holds consistently across major backbone models and question types, with performance gains ranging from 2%–86% over vanilla baselines and 2%–88% over few-shot baselines. However, RAG exhibits a precision-recall tradeoff for complex reasoning MAQs, achieving higher precision but lower recall across all main models. This suggests benchmarks should explicitly account for task priorities (precision-oriented vs. recall-oriented scenarios) when evaluating QA systems.
@@
While few-shot learning marginally improves binary and MCQ responses by clarifying answer formats, 33% of its implementations underperform vanilla prompts. All main models except GPT-4 show metric degradation in $\geq$2 categories, with MAQs being most vulnerable: LLMs often misinterpret few-shot examples as definitive answer templates. Appendix A data reveals extreme cases (e.g., Vicuna-13B’s 90% metric decline), highlighting the risk of overfitting to limited examples. This implies that few-shot strategies require careful validation against answer diversity in target domains.
5.3 Supervised Fine-tuning on XUBench

### v3

# Benchmarking for Domain-Specific LLMs: A Case Study on Academia and Beyond

@@ Abstract
The increasing demand for domain-specific evaluation of large language models (LLMs) has led to the development of numerous benchmarks. These efforts often adhere to the principle of data scaling, relying on large corpora or extensive question-answer (QA) sets to ensure broad coverage. However, the impact of corpus and QA set design on the precision and recall of domain-specific LLM performance remains poorly understood. In this paper, we argue that data scaling is not always the optimal principle for domain-specific benchmark construction. Instead, we introduce Comp-Comp, an iterative benchmarking framework grounded in the principle of comprehensiveness and compactness. Comprehensiveness ensures semantic recall by covering the full breadth of the domain, while compactness improves precision by reducing redundancy and noise. To demonstrate the effectiveness of our approach, we present a case study conducted at a well-renowned university, resulting in the creation of PolyBench, a large-scale, high-quality academic benchmark. Although this study focuses on academia, the Comp-Comp framework is domain-agnostic and readily adaptable to a wide range of specialized fields. The source code and datasets can be accessed at https://github.com/Anya-RB-Chen/COMP-COMP.
Benchmarking for Domain-Specific LLMs: A Case Study on
Academia and Beyond
Rubing Chen1, Jiaxin Wu1, Jian Wang1, Xulu Zhang1, Wenqi Fan1, Chenghua Lin2, Xiao-Yong Wei1, thanks: Corresponding author, Qing Li1 1The Hong Kong Polytechnic University  2The University of Manchester
@@ 1 Introduction
Figure 1: Illustration of comprehensiveness and compactness in benchmarking. (a) Comprehensiveness ensures broad semantic coverage of the domain, while compactness improves semantic precision by minimizing redundancy, leading to a more uniform data distribution in domain corpora. (b) Semantic distributions of representative benchmarks in the medical domain.
@@
Large language models (LLMs) have demonstrated impressive capabilities across a wide range of tasks OpenAI (2024). While general-purpose LLMs excel in broad applications, domain-specific LLMs have emerged as critical tools for delivering precise and accurate responses tailored to specific user needs Fei et al. (2023); Cai et al. (2024); Chang et al. (2024). These models are particularly valuable in fields where fixed-answer questions are prevalent, such as law, medicine, and academic, where precision is paramount Ge et al. (2024). However, the development of domain-specific LLMs necessitates high-quality benchmarks that can effectively evaluate their performance. Such benchmarks must not only ensure accurate responses within the domain, but also maintain broad coverage of the domain’s semantic space to avoid catastrophic forgetting, where models lose proficiency in broader contexts Luo et al. (2023). This dual requirement underscores the need to develop effective approaches for domain-specific benchmarking.
Benchmark Domain Closeness QA Type # Questions
SciEval Cai et al. (2024) Science Open MCQ, QA 16,522
DISC-Law-Eval Yue et al. (2023) Law Open MCQ, MAQ, QA 2,863
Counseling Bench Liu et al. (2023) Medical Open QA 229
CMtMedQA Yang et al. (2024) Medical Open QA 70,000
COVID-QA Möller et al. (2020) Healthcare Closed QA 2,019
WC2014QA Zhang et al. (2016) Sports Closed QA 10,211
PolyBench (Ours) Academic Closed Binary, MAQ, QA, MCQ 24,883
Table 1: Comparison between domain-specific benchmarks.
Previous studies on constructing domain-specific benchmarks have primarily relied on the principle of data scaling, utilizing extensive corpora or large question sets to ensure broad coverage Cai et al. (2024); Fei et al. (2023). While this approach is widespread and forms the basis for many benchmark designs, it presents several limitations that can affect the efficacy of model evaluation. First, many benchmarks are derived from skill examinations, practice exercises, or expert-curated datasets, often overlooking a critical assessment of their effectiveness Liu et al. (2023); Ling et al. (2023). Although these benchmarks are diverse in question-answering types, ranging from multiple-choice to open-ended formats, they fail to systematically examine the semantic alignment between questions and the underlying corpus Yue et al. (2023); Cui et al. (2023). This lack of alignment means that the benchmarks may not fully represent the domain’s knowledge structure, leading to evaluations that do not accurately reflect a model’s true capabilities within the specific domain. Second, domain-specific benchmarks often employ LLMs for scoring but neglect to ensure consistency between training and evaluation datasets Zhang et al. (2023); Yang et al. (2024). This oversight can introduce biases where models inadvertently perform better on evaluation tasks due to similarities with the training data, rather than a genuine understanding of the domain. It highlights the need for a more balanced approach to benchmarking, one that integrates both horizontal coverage across a wide range of topics and vertical depth within specific subdomains.
@@
In this paper, we propose Comprehensiveness-Compactness (Comp-Comp), a novel framework for constructing domain-specific benchmarks that dynamically balances the semantic distribution of corpora and benchmark questions. As illustrated in Figure 1, Comp-Comp iteratively expands the range of corpora and questions, continuously assessing their comprehensiveness (maximizing recall) and compactness (optimizing precision). The framework introduces three crucial mechanisms: (1) Comprehensiveness and Compactness Monitoring: We encode the corpora and questions into a unified semantic space, using Gaussian Kernel Density Estimation (KDE) to improve these metrics iteratively Terrell and Scott (1992). (2) Corpora-Question Interaction: We ensure comprehensive evaluation by generating new questions from underrepresented areas in the corpora. (3) Superiority in Closed Domains: We prioritize achieving a balance between comprehensiveness and compactness, ensuring the framework excels in constructing closed-domain benchmarks that demand precise corpus evaluation and accurate answer verification.
@@
To validate our framework, we conducted a case study in a well-renowned university and applied Comp-Comp to develop PolyBench, one of the first and most extensive benchmarks for the academic domain (see Table 1). PolyBench includes 24.9k questions across various formats and incorporates user-interest-oriented forum discussions, setting a new standard for domain-specific benchmarking. Extensive experiments on the PolyBench demonstrate the practicality and effectiveness of Comp-Comp, providing a scalable approach for future benchmarking in other domains.
@@
In summary, our contributions are as follows: (1) We propose Comp-Comp, a principled approach that balances comprehensiveness and compactness in domain-specific benchmark construction. (2) We develop PolyBench, a large-scale, high-quality benchmark for the academic domain, which addresses the gap of lacking well-established academic LLM benchmarks. (3) Our method is extensible across different domains, offering valuable insights for future benchmarking. By addressing the limitations of existing methods and providing a systematic benchmark design, our work advances the development of domain-specific LLMs.
@@ 2 Related Work
Domain-specific evaluation has seen significant advances across critical verticals, each adopting distinct strategies reflective of their unique demands. In the legal domain, systems like DISC-LawLLM Yue et al. (2023) employ a dual evaluation paradigm: standardized law examinations (MCQs with tiered difficulty) paired with ChatGPT, arbitrated subjective assessments. LawBench Fei et al. (2023) further refines this by grounding evaluations in Bloom’s Taxonomy Krathwohl (2002), systematically testing knowledge memorization, understanding, and application within Chinese legal contexts. The medical domain presents alternative approaches: ChatCounselor Liu et al. (2023) evaluates seven mental consultation dimensions via GPT-4, while CMtMedQA Yang et al. (2024) uses dialogue-based capacity assessments scored across three dimensions. Even within single domains, fragmentation persists: Figure 1(b) reveals disjointed semantic distributions across medical benchmarks, including MedExQA Kim et al. (2024), MMLU Hendrycks et al. (2021), PubMedQA Jin et al. (2019), and MedQA Jin et al. (2021), with clear boundaries between corpora despite shared topical grounding. Educational benchmarks like C-Eval Huang et al. (2024), though comprehensive, exhibit similar fragmentation, as noted in EduChat’s evaluation Dan et al. (2023). These efforts collectively highlight a critical gap: the absence of unified metrics to assess whether benchmarks sufficiently cover a domain’s semantic space.
@@ 3.1 Comprehensiveness and Compactness Assessment
Given a set of data point $X=\{x_{1},x_{2},...,x_{m}\}$, its comprehensiveness is measured by computing its coverage of a space $S$ with data points $\{s_{1},s_{2},...,s_{n}\}$. To achieve this, we encode all data points in $X$ and $S$ to the semantic space as follows:
@@
On the other hand, the compactness of the corpora and QA is controlled by whether the inclusion of new data points $X$ gives much semantic residency to the existing set $Y$. Thus, we measure the Pearson correlation coefficient Cohen et al. (2009) between the semantic distributions of $X$ and $Y$ as:
@@ 3.3 Mining User Interests for QA Generation
Moreover, in the QA generation, we propose that we should not only generate QAs based on the corpora $C$ but also include user-interested questions. For example, the frequently asked questions and the questions discussed in public forums (e.g., Quora and Zhihu). To this end, we add user-interested questions as a special gap points to $A_{gap}^{\prime}(C,Q)$.
4 PolyBench: An Academic Benchmark
In this section, we employ the proposed framework to construct PolyBench, a comprehensive and compact benchmark in the academic domain.
@@ 4.1 Assessment Guided Corpora Collection
Table 2: Statistics of the PolyBench.
@@ 5 Experiments
Table 3: The performance of In-Context Learning (ICL) and Supervised Fine-tuning (SFT) approaches (i.e., few-shot learning and RAG) on the PolyBench with various types of questions.
@@ 5.1 Experimental Settings
We include both frozen and supervised fine-tuning (SFT) LLMs with in-context learning (ICL) prompts as baselines. For the ICL, we apply few-shot learning Parnami and Lee (2022) and Retrieval Augmented Generation (RAG) Lewis et al. (2020); Gao et al. (2023) to include domain knowledge in the prompt for LLMs. For the SFT, we use LoRA Hu et al. (2021) to fine-tune several LLM backbones. For LLM backones, we apply both close-sourced and open-sourced LLMs including GPT3.5, GPT4 OpenAI (2024), Vicuna-7B, Vicuna-13B Chiang et al. (2023), LLaMA3-8B Touvron et al. (2023), GLM4-9B GLM et al. (2024), and Qwen2-7B Bai et al. (2023).
@@
For the PolyBench construction, we set the key hyperparameters based on preliminary experiments aimed at balancing comprehensiveness and compactness. The corpus compactness threshold was set to $t_{c}=0.05$ in Eq. (7) for the corpora collection and $h=5.0$ in Eq.(3) and Eq. (4), and $t_{d}=0.6$ in Eq (8) for the QA generation. Results of more parameters and their sensitivity will be presented in Section 5.5. We use Gte-Large Li et al. (2023) encoder to encode data chunks, selected for its robust performance in capturing nuanced semantic similarities.
5.2 In-Context Learning on PolyBench
Table 3 compares the performances of different in-context-learning techniques, i.e., few-shot learning and RAG, on answering several different QA sets of PolyBench. Experiment results on more LLMs are listed in Appendix A’s Table 7. Based on the results, we have several findings:
@@ 5.2 In-Context Learning on PolyBench
Our analysis demonstrates that RAG substantially improves precision and recall over vanilla and few-shot learning in domain-specific QA. This holds consistently across major backbone models and question types, with performance gains ranging from 2%–86% over vanilla baselines and 2%–88% over few-shot baselines. However, RAG exhibits a precision-recall tradeoff for complex reasoning MAQs, achieving higher precision but lower recall across all main models. This suggests benchmarks should explicitly account for task priorities (precision-oriented vs. recall-oriented scenarios) when evaluating QA systems.
@@
While few-shot learning marginally improves binary and MCQ responses by clarifying answer formats, 33% of its implementations underperform vanilla prompts. All main models except GPT-4 show metric degradation in $\geq$2 categories, with MAQs being most vulnerable: LLMs often misinterpret few-shot examples as definitive answer templates. Appendix A data reveals extreme cases (e.g., Vicuna-13B’s 90% metric decline), highlighting the risk of overfitting to limited examples. This implies that few-shot strategies require careful validation against answer diversity.
5.3 Supervised Fine-tuning on PolyBench
@@ 5.4 Comparison between ICL and SFT
Method Helpfulness Conciseness Correctness BLEU-2 | -4
LLaMA3-8B
Vanilla 3.70±0.99 3.67±0.89 3.53±1.02 0.19 | 0.13
Fewshot 3.73±0.88 3.66±0.95 3.58±1.02 0.23 | 0.15
RAG 3.11±0.96 3.32±0.99 3.27±1.11 0.35 | 0.30
GLM4-9B
Vanilla 3.64±0.97 3.62±0.99 3.58±0.92 0.19 | 0.12
Fewshot 3.60±1.02 3.62±1.00 3.51±0.95 0.20 | 0.13
RAG 3.19±0.84 3.07±0.83 3.30±0.96 0.31 | 0.26
GPT-4
Vanilla 3.12±0.79 2.97±0.78 3.12±0.84 0.29 | 0.20
Fewshot 2.75±0.72 3.06±0.89 2.86±0.87 0.31 | 0.22
RAG 3.11±0.77 2.97±0.88 3.14±0.94 0.36 | 0.32
LLaMA3-8B (FT)
Vanilla 2.41±0.92 2.63±0.91 2.62±1.00 0.28 | 0.19
Fewshot 2.40±0.93 2.77±1.09 2.56±0.99 0.29 | 0.20
RAG 2.58±0.99 2.85±1.09 2.74±1.09 0.43 | 0.38
GLM4-9B (FT)
Vanilla 2.93±1.06 3.03±1.02 2.84±1.16 0.28 | 0.20
Fewshot 2.89±0.96 3.08±1.03 2.99±1.03 0.29 | 0.20
RAG 2.95±0.94 3.06±0.99 2.99±1.10 0.44 | 0.40
Qwen2-7B (FT)
Vanilla 2.99±1.03 3.08±1.16 2.99±1.20 0.26 | 0.18
Fewshot 3.03±1.10 3.07±1.11 2.96±1.22 0.26 | 0.18
RAG 2.78±1.15 2.95±1.17 2.92±1.23 0.40 | 0.35
Table 5: User Survey Results and Open-Ended BLEU Scores of PolyBench (mean ± standard deviation).
Dataset $t_{d}$ thresholds for questions $Q$ (zero-shot) $t_{c}$ thresholds for corpora $C$ (RAG)
w/o $t_{d}$ 0.1 0.3 0.5 0.7 0.9 w/o $t_{c}$ 0.0 0.05 0.1 0.15 0.2 0.25 0.3
MedQA # $Q$ 1273 21 24 28 36 503 # $C$ 213330 21342 85338 96004 117336 181332 181332 202664
Acc. 0.684 0.714 0.750 0.786 0.778 0.660 Acc. 0.870 0.874 0.855 0.850 0.846 0.866 0.874 0.854
PubMedQA # $Q$ 1000 14 29 35 92 347 # $C$ 3358 2356 3358 3358 3358 3358 3358 3358
Acc. 0.462 0.429 0.586 0.543 0.467 0.467 Acc. 0.670 0.681 0.670 0.670 0.670 0.670 0.670 0.670
FiQA # $Q$ 1706 117 287 426 528 626 # $C$ 15404 2310 7704 13094 14634 15404 15404 15404
BLEU-2 0.046 0.042 0.045 0.047 0.047 0.046 BLEU-2 0.054 0.051 0.051 0.052 0.055 0.054 0.054 0.054
FinanceBench # $Q$ 150 11 32 48 70 85 # $C$ 180 54 108 162 180 180 180 180
Acc. 0.007 0.000 0.000 0.000 0.014 0.012 Acc. 0.053 0.007 0.007 0.053 0.053 0.053 0.053 0.053
Table 6: Performance of the zero-shot QA ($t_{d}$ thresholds for questions $Q$) and RAG ($t_{c}$ thresholds for corpora $C$) baseline on different datasets across varying thresholds $t_{d}$ and $t_{c}$, respectively.